OpenStudy (j3nny0rtiz):

How should this problem be solved? (a3)2 A. a3 + 2 = a5 B. a3 ÷ 2 = a1.1 C. a3 – 2 = a1 D. a3 x 2 = a6

OpenStudy (anonymous):

Pretty sure D

OpenStudy (j3nny0rtiz):

thank u

OpenStudy (anonymous):

\[\left(a^3\right)^2=a^{3\ 2}=a^6 \]